# Creativity in AI: Progresses and Challenges

### v2

@@ 1. Introduction
While these remarkable achievements can be seen as signs of the presence of creative capacity in transformer-based language models, it should be noted that these models rely on an astronomically large number of parameters and are trained on massive amounts of public and private data (Brown et al., 2020a). Hence, it is not entirely clear whether the seemingly extraordinary outputs of these models are the result of a truly creative inner process and robust generalization or the result of powerful interpolation and strong memorization skills (McCoy et al., 2023; Hupkes et al., 2022; Carlini et al., 2022; Bender et al., 2021; Marcus, 2020; Bender and Koller, 2020). Recent works have shown that language models fail at real-world commonsense reasoning and compositionality tasks (Ismayilzada et al., 2023; Dziri et al., 2023), occasionally copy large amounts of text from their training data (McCoy et al., 2021), significantly lag behind experts in creative writing (Chakrabarty et al., 2023a), produce less diverse content (Anderson et al., 2024a; Padmakumar and He, 2023), struggle with creative problem-solving (Tian et al., 2023) and abstract reasoning (Gendron et al., 2023; Mitchell et al., 2023) and suffer from factual inconsistency and hallucination issues (Banerjee et al., 2024; Elazar et al., 2021). While previous works have reviewed developments on some aspects of AI creativity (Amin and Burghardt, 2020; Lai and Nissim, 2024; Oliveira, 2017; Elzohbi and Zhao, 2023; Franceschelli and Musolesi, 2021b; Rowe and Partridge, 1993), a more holistic and broader review of the field is necessary to understand its rapid advancements.
@@
In this work, we provide the general AI audience with a timely summary of the state of the creative capabilities of the latest AI systems. While creativity is a broad concept that can be explored in a wide range of areas, in this survey, we focus on four main areas where machine creativity has been extensively investigated: Linguistic Creativity (§3), Creative Problem-Solving (§4), Artistic Creativity (§5) and Scientific Creativity (§6). These areas capture four crucial pillars of creative thinking in humans: linguistic creativity enables us to manipulate language in novel ways for effective communication of ideas; creative problem-solving helps us find efficient solutions by thinking out-of-the-box; artistic creativity allows for the expression of emotions, ideas, and aesthetics through various media; and scientific creativity drives innovation by enabling the formulation of new hypotheses, theories, and discoveries. Together, these areas represent distinct yet interconnected facets of creativity, providing a comprehensive framework for studying how machines can emulate or assist human-like creative processes across different domains. For each area, we survey representative tasks, resources, methods, and major findings and present a taxonomy of these works in Figure 2. Our review indicates that although the latest AI models are generally proficient in generating linguistically and artistically creative outputs, such as poems, images, and music, they face challenges with tasks demanding creative problem-solving, abstract reasoning, and compositionality. Their outputs often lack diversity and originality, exhibit long-term incoherence, and are prone to hallucinations. We also briefly discuss the emerging challenges brought by generative models concerning copyright and authorship of artworks (§7). Finally, in the last section (§8), we argue for a comprehensive evaluation of creativity in AI that considers several dimensions of creativity and the creative process at its core. Furthermore, we discuss future research directions to enhance the creativity of AI systems, potentially drawing ideas from cognitive science and psychology.
@@ 2.1. Definitions
The novelty criterion is typically self-explanatory to the point that people equate it to creativity in everyday life. This is generally known as intrapersonal or personal creativity (often denoted as P-creativity), i.e. the product is novel within the frame of a person’s life (Boden, 2004; Stein, 1953; Weisberg, 1986). Researchers distinguish it from the interpersonal or historical creativity (often denoted as H-creativity), i.e. the product is novel with respect to the entire history of people such as Einstein’s general relativity theory.
Many theorists have argued that novelty is insufficient for creativity, and value dimension is needed to filter out original nonsense, such as something generated by a truly random process. While value is generally understood as something inherently “good” for the respective audience, there appears to be such a thing as malevolent or “dark” creativity. For instance, one can be creative in producing torture instruments or in committing terrorist atrocities (Gaut, 2010). Therefore, the interpretation of the value of a product as being “effective” towards its intended end, regardless of whether that end is morally good or bad, has been suggested as a better alternative (Livingston, 2018). However, we should note that evaluating utility or value still requires an outside judgment which is subjective, can be faulty or biased and can change over time and across cultures. This is especially apparent in arts as there are many great artists (Bach, Van Gogh etc.) whose “value” have only been recognized longtime after their death. Moreover, sometimes novelty is itself the value created by the artist because no one has done it before, particularly, in visual arts. Hence, some researchers have recently argued to drop the value criterion altogether from the definition of creativity (Brandt, 2021; Weisberg, 2015b).
Despite its wide adoption, the sufficiency of the novelty and value conditions for creativity has also been challenged (Gaut, 2010; Arnheim, 2001; Weisberg, 2015a). It has been argued that agency, a capacity to have beliefs, desires or intentional states, is a required attribute of creativity. For example, Gaut (2010) mentions the tectonic movement of the earth’s crust that can produce valuable (financially and aesthetically) and sometimes original (new variation) diamonds, but we would hardly call tectonic movements creative. However, mere agency without intentionality is also insufficient. Gaut (2010) illustrates this with an example where a person walking in a studio accidentally knocks over a set of paints, which spill onto a canvas and happen to create a beautiful and original painting. Weisberg (2015b) has gone even further to suggest that creativity is simply intentional novelty.
@@
2.2. Types
@@
While creativity manifests itself in various forms across domains, even within a particular domain, different types of creativity can be distinguished based on the difficulty level of the inherent creative process involved. Four-C model of creativity is one such popular theory that differentiates between four types of creativity corresponding to four levels of difficulty involved in producing creative artifacts (Kaufman and Beghetto, 2009). The first major type of creativity is known as little-c creativity which is what we find in everyday life as solutions to minor problems. Examples might include combining unusual ingredients to make a new type of meal or using a hand-held vacuum cleaner on the ceiling to remove flies. Almost everyone possesses this type of creativity in one way or another. The second main type of creativity in this model is the Big-C creativity that includes major works of scientific, technological, social, or artistic importance. Examples could be Darwin’s theory of evolution, the invention of the printing press, or Leonardo Da Vinci’s painting of the Mona Lisa. In addition to these two major categories, Kaufman and Beghetto (2009) also defines two minor categories of creativity. First is the mini-c creativity for very small-scale cases of creativity such as young children’s drawing or their creative experiments with Lego pieces. Second is the Pro-C creativity which is proposed for work produced by professional but non-prominent practitioners such as professional musicians or artists who generate novel work, but do not make historical contributions.
@@
Recently, there has been a suggestion to differentiate three types of creativity corresponding to three major levels of innovation that can be achieved (Hassabis, 2018). First can be achieved through interpolation where a prototypical creative artifact is produced by averaging all the artifacts of the same class seen before. While it is an original product that did not exist before, it still relies heavily on the other existing products. An example would be to come up with a novel winning strategy in chess that is a combination of existing different strategies. Consequently, a second type of innovation can be achieved through extrapolation where a creative artifact extends the boundaries of what has been seen before, but is still of the same class. A completely new chess move that is not related to any existing moves can be seen as an example of extrapolative creativity. Finally, the highest level of creativity can be termed as invention where a creative artifact introduces a novel class of its own. Inventing chess itself or any major scientific invention is a perfect example of this type of creativity.
@@ 2.3.1. Manual Evaluation
Since creative products vary greatly in their forms and are hard to characterize with objective measures, the simplest and most common way to evaluate them is to ask other humans to manually rate them based on some criteria associated with creativity, which differs from task to task (Lamb et al., 2015). For example, in story generation, humans are typically asked to rate a generated story on aspects such as interestingness, coherence, relevance, humanlikeness and etc. (Goldfarb-Tarrant et al., 2020; Rashkin et al., 2020; Yang et al., 2022). In other tasks where the goal is to produce multiple responses such as common psychometric creativity tests Alternative Uses Task (AUT) (Guilford, 1967) and Torrance Tests of Creative Thinking (TTCT) (Torrance, 1974), evaluation is centered around four dimensions of creativity: fluency (the total number of meaningful, and relevant ideas generated in response to the stimulus), flexibility (the number of different categories of relevant responses), originality (the uniqueness or rarity of responses) and elaboration (the amount of detail in the responses).
@@ 2.3.2. Automated Evaluation
This dimension is generally the hardest to evaluate as it depends on the subjective utility or performance of the artifact which is typically judged by domain experts of that artifact and can radically change across domains (Maher, 2010). In visual arts, this might correspond to “beauty”, whereas in science to “logical correctness”. Therefore, a metric appropriate for its domain should be employed. For example, in open-ended story generation, a minimally useful story can be defined as a relevant, coherent, and meaningful story. In this sense, automated metrics measuring the overall quality of a story can be leveraged (Xie et al., 2023a; Guan and Huang, 2020; Chen et al., 2022b), however, it is often challenging to measure coherence (Zhao et al., 2023b; Laban et al., 2021). Another more general evaluation of utility has been suggested by Franceschelli and Musolesi (2022) based on the discriminator score in GANs.
@@ Methods
Humorous content (Ravi et al., 2024; Hessel et al., 2023b; Xie et al., 2023b; Arora et al., 2022; Peyrard et al., 2021; Kayatani et al., 2021),
@@
Morphological Productivity (Ismayilzada et al., 2024; Anh et al., 2024; Weissweiler et al., 2023)
@@
3. Linguistic Creativity
@@
The creative aspect of language in linguistics has been discussed since the early days (Chomsky, 1965). Chomsky, in this paper, attributes creativity mainly to the essential property of language to provide means to express many thoughts indefinitely. However, several linguists since Chomsky have argued against using this characterization since it does not align with the everyday definition of creativity (Bergs, 2019; Sampson, 2017; Zawada, 2006). Chomsky’s theory of grammar might generate an infinite number of sentences; it, however, relies on a fixed set of rules, while creativity requires deviation from rules. In this sense, Sampson (2017) suggests distinguishing between F-creativity (fixed) and E-creativity (extending), where F-creativity refers to the Chomskian interpretation of linguistic creativity (a.k.a productivity in morphology) and E-creativity corresponds to the real linguistic innovation such as metaphors, jokes, neologisms, etc. Some recent works have explored the F-creativity of large language models and found that this task is challenging in general and even harder in more morphological complex languages (Ismayilzada et al., 2024; Anh et al., 2024; Weissweiler et al., 2023). Most past works however have focused on studying the E-creativity of AI systems which we review in the following sections.
3.1. Humor
@@
Humor is one of the most common ways humans creatively use language to express their ideas and feelings. Early works to model humor focused on hand-crafted linguistic templates and wordplay (Stock and Strapparava, 2005; Taylor and Mazlack, 2004; Raskin and Attardo, 1994). Subsequent works have leveraged language’s lexical and syntactic properties as humor-specific features for humor detection (Liu et al., 2018b; Yang et al., 2015; Zhang and Liu, 2014). The growing interest in computational humor in recent years has resulted in several shared tasks organized by the NLP community (Meaney et al., 2021; Hossain et al., 2020a; Van Hee et al., 2018; Castro et al., 2018; Potash et al., 2017; Miller et al., 2017). Latest works have developed methods based on neural networks and language models to generate and detect humorous content (Ravi et al., 2024; Arora et al., 2022; Peyrard et al., 2021; Annamoradnejad and Zoghi, 2020; Ziser et al., 2020; Amin and Burghardt, 2020; Hossain et al., 2019; Chen and Soo, 2018; Bertero and Fung, 2016), jokes (Horvitz et al., 2024; Tang et al., 2022; Xie et al., 2021; Weller and Seppi, 2019; Ren and Yang, 2017), puns (He et al., 2019; Mittal et al., 2022; Yu et al., 2018), and sarcasm (Chakrabarty et al., 2022c, 2020a). Several datasets have also been proposed to benchmark the humor capacity of LLMs in several languages including English (Horvitz et al., 2024; Jain et al., 2024; Tang et al., 2022; Meaney et al., 2021; Hossain et al., 2020b, 2019; Miller et al., 2017), Chinese (Zhang et al., 2019b), Italian (Buscaldi and Rosso, 2007), Spanish (Castro et al., 2017), Dutch (Winters and Delobelle, 2020) and Russian (Blinov et al., 2019). Computational humor has also been explored in multimodal settings involving images, audio, and video in addition to text (Hessel et al., 2023a; Xie et al., 2023b; Hasan et al., 2019; Bertero and Fung, 2016; Radev et al., 2016; Shahaf et al., 2015). While the latest methods particularly LLMs show an impressive ability to generate and detect humorous content, recent work has also shown that these models still fail to reliably understand humor (Hessel et al., 2023a; Góes et al., 2023a; Kocoń et al., 2023; Borji, 2023a) and generated jokes typically lack diversity (Jentzsch and Kersting, 2023) which has been attributed to training on less diverse humor datasets (Baranov et al., 2023). Creative training frameworks have also been developed to improve the humor generation capabilities of LLMs (Zhong et al., 2023). We refer the reader to Amin and Burghardt (2020) for an in-depth survey on computational humor.
3.2. Figurative Language

### v3

@@ 1. Introduction
While these remarkable achievements can be seen as signs of the presence of creative capacity in transformer-based language models, it should be noted that these models rely on an astronomically large number of parameters and are trained on massive amounts of public and private data (Brown et al., 2020a). Hence, it is not entirely clear whether the seemingly extraordinary outputs of these models are the result of a truly creative inner process and robust generalization or the result of powerful interpolation and strong memorization skills (McCoy et al., 2023; Hupkes et al., 2022; Carlini et al., 2022; Bender et al., 2021; Marcus, 2020; Bender and Koller, 2020). Recent works have shown that language models fail at real-world commonsense reasoning and compositionality tasks (Ismayilzada et al., 2023; Dziri et al., 2023), occasionally copy large amounts of text from their training data (Lu et al., 2024b; McCoy et al., 2021), significantly lag behind humans in creative writing (Ismayilzada et al., 2024b; Chakrabarty et al., 2023a), produce less diverse content (Anderson et al., 2024a; Padmakumar and He, 2023), struggle with creative problem-solving (Huang et al., 2024b; Tian et al., 2023) and abstract reasoning (Gendron et al., 2023; Mitchell et al., 2023) and suffer from factual inconsistency and hallucination issues (Banerjee et al., 2024; Elazar et al., 2021). While previous works have reviewed developments on some aspects of AI creativity (Amin and Burghardt, 2020; Lai and Nissim, 2024; Oliveira, 2017; Elzohbi and Zhao, 2023; Franceschelli and Musolesi, 2021b; Rowe and Partridge, 1993), a more holistic and broader review of the field is necessary to understand its rapid advancements.
@@
In this work, we provide the general AI audience with a timely summary of the state of the creative capabilities of the latest AI systems. While creativity is a broad concept that can be explored in a wide range of areas, in this survey, we focus on four main areas where machine creativity has been extensively investigated: Linguistic Creativity (§3.1), Creative Problem-Solving (§3.2), Artistic Creativity (§3.3) and Scientific Creativity (§3.4). These areas capture four crucial pillars of creative thinking in humans: linguistic creativity enables us to manipulate language in novel ways for effective communication of ideas; creative problem-solving helps us find efficient solutions by thinking out-of-the-box; artistic creativity allows for the expression of emotions, ideas, and aesthetics through various media; and scientific creativity drives innovation by enabling the formulation of new hypotheses, theories, and discoveries. Together, these areas represent distinct yet interconnected facets of creativity, providing a comprehensive framework for studying how machines can emulate or assist human-like creative processes across different domains. For each area, we survey representative tasks, resources, methods, and major findings and present a taxonomy of these works in Figure 2. Our review indicates that although the latest AI models are generally proficient in generating linguistically and artistically creative outputs, such as poems, images, and music, they face challenges with tasks demanding creative problem-solving, abstract reasoning, and compositionality. Their outputs often lack diversity and originality, exhibit long-term incoherence, and are prone to hallucinations. We also briefly discuss the emerging challenges brought by generative models concerning copyright and authorship of artworks (§4). Finally, in the last section (§5), we argue for a comprehensive evaluation of creativity in AI that considers several dimensions of creativity and the creative process at its core. Furthermore, we discuss future research directions to enhance the creativity of AI systems, potentially drawing ideas from cognitive science and psychology.
@@ 2.1. Defining Creativity
The novelty criterion is typically self-explanatory to the point that people equate it to creativity in everyday life. However, many theorists have argued that novelty is insufficient for creativity, and value dimension is needed to filter out original nonsense, such as something generated by a truly random process. While value is generally understood as something inherently “good” for the respective audience, there appears to be such a thing as malevolent or “dark” creativity. For instance, one can be creative in producing torture instruments or in committing terrorist atrocities (Gaut, 2010). Therefore, the interpretation of the value of a product as being “effective” towards its intended end, regardless of whether that end is morally good or bad, has been suggested as a better alternative (Livingston, 2018). However, we should note that evaluating utility or value still requires an outside judgment which is subjective, can be faulty or biased and can change over time and across cultures. This is especially apparent in arts as there are many great artists (Bach, Van Gogh etc.) whose “value” have only been recognized longtime after their death. Moreover, sometimes novelty is itself the value created by the artist because no one has done it before, particularly, in visual arts. Hence, some researchers have recently argued to drop the value criterion altogether from the definition of creativity (Brandt, 2021; Weisberg, 2015a).
Despite its wide adoption, the sufficiency of the novelty and value conditions for creativity has also been challenged (Gaut, 2010; Arnheim, 2001; Weisberg, 2015b). It has been argued that agency, a capacity to have beliefs, desires or intentional states, is a required attribute of creativity. For example, Gaut (2010) mentions the tectonic movement of the earth’s crust that can produce valuable (financially and aesthetically) and sometimes original (new variation) diamonds, but we would hardly call tectonic movements creative. However, mere agency without intentionality is also insufficient. Gaut (2010) illustrates this with an example where a person walking in a studio accidentally knocks over a set of paints, which spill onto a canvas and happen to create a beautiful and original painting. Weisberg (2015a) has gone even further to suggest that creativity is simply intentional novelty.
@@
2.2. Types of Creativity
While creativity manifests itself in various forms across domains, even within a particular domain, different types of creativity can be distinguished based on its timing or target audience and the difficulty level of the inherent creative process involved. A person might come up with a creative idea that is new to him/her but already invented by someone else in history. This is generally known as intrapersonal or personal (a.k.a psychological) creativity (often denoted as P-creativity), i.e. the product is novel within the frame of a person’s life (Boden, 2004; Stein, 1953; Weisberg, 1986). Researchers distinguish it from the interpersonal or historical creativity (often denoted as H-creativity), i.e. the product is novel with respect to the entire history of people such as Einstein’s general relativity theory.
@@ 2.2. Types of Creativity
Four-C model of creativity, on the other hand, differentiates between four types of creativity corresponding to four levels of difficulty involved in producing creative artifacts (Kaufman and Beghetto, 2009). The first major type of creativity is known as little-c creativity which is what we find in everyday life as solutions to minor problems. Examples might include combining unusual ingredients to make a new type of meal or using a hand-held vacuum cleaner on the ceiling to remove flies. Almost everyone possesses this type of creativity in one way or another. The second main type of creativity in this model is the Big-C creativity that includes major works of scientific, technological, social, or artistic importance. Examples could be Darwin’s theory of evolution, the invention of the printing press, or Leonardo Da Vinci’s painting of the Mona Lisa. In addition to these two major categories, Kaufman and Beghetto (2009) also defines two minor categories of creativity. First is the mini-c creativity for very small-scale cases of creativity such as young children’s drawing or their creative experiments with Lego pieces. Second is the Pro-C creativity which is proposed for work produced by professional but non-prominent practitioners such as professional musicians or artists who generate novel work, but do not make historical contributions.
@@
Recently, there has been a suggestion to differentiate three types of creativity corresponding to three major levels of innovation that can be achieved (Hassabis, 2018). First can be achieved through interpolation where a prototypical creative artifact is produced by averaging all the artifacts of the same class seen before. While it is an original product that did not exist before, it still relies heavily on the other existing products. An example would be to come up with a novel winning strategy in chess that is a combination of existing different strategies. Consequently, a second type of innovation can be achieved through extrapolation where a creative artifact extends the boundaries of what has been seen before, but is still of the same class. A completely new chess move that is not related to any existing moves can be seen as an example of extrapolative creativity. Finally, the highest level of creativity can be termed as invention where a creative artifact introduces a novel class of its own. Inventing chess itself or any major scientific invention is a perfect example of this type of creativity. This type of creativity typically requires transformation of the existing conceptual space and is also known as transformational creativity (Boden, 2004).
@@ 2.3.1. Manual Evaluation
Since creative products vary greatly in their forms and are hard to characterize with objective measures, the simplest and most common way to evaluate them is to ask other humans to manually rate them based on some criteria associated with creativity, which differs from task to task (Lamb et al., 2015). For example, in story generation, humans are typically asked to rate a generated story on aspects such as interestingness, coherence, relevance, humanlikeness and etc. (Yang and Jin, 2024; Goldfarb-Tarrant et al., 2020; Rashkin et al., 2020; Yang et al., 2022). In other tasks where the goal is to produce multiple responses such as common psychometric creativity tests Alternative Uses Task (AUT) (Guilford, 1967) and Torrance Tests of Creative Thinking (TTCT) (Torrance, 1974), evaluation is centered around four dimensions of creativity: fluency (the total number of meaningful, and relevant ideas generated in response to the stimulus), flexibility (the number of different categories of relevant responses), originality (the uniqueness or rarity of responses) and elaboration (the amount of detail in the responses).
@@ 2.3.2. Automated Evaluation
This dimension is generally the hardest to evaluate as it depends on the subjective utility or performance of the artifact which is typically judged by domain experts of that artifact and can radically change across domains (Maher, 2010). In visual arts, this might correspond to “beauty”, whereas in science to “logical correctness”. Therefore, a metric appropriate for its domain should be employed. For example, in open-ended story generation, a minimally useful story can be defined as a relevant, coherent, and meaningful story. In this sense, automated metrics measuring the overall quality of a story can be leveraged (Xie et al., 2023a, b; Guan and Huang, 2020; Chen et al., 2022b), however, it is often challenging to measure coherence (Zhao et al., 2023b; Laban et al., 2021). Another more general evaluation of utility has been suggested by Franceschelli and Musolesi (2022) based on the discriminator score in GANs. Since in GANs, the discriminator learns the distribution of the real (and valuable) data, its score can directly be used as a proxy metric to measure value.
@@ Methods
Humorous content (Ravi et al., 2024; Hessel et al., 2023b; Xie et al., 2023c; Arora et al., 2022; Peyrard et al., 2021; Kayatani et al., 2021),
@@
Morphological Productivity (Ismayilzada et al., 2024a; Anh et al., 2024; Weissweiler et al., 2023)
@@
3. Domains of Creativity
Creativity is a multifaceted concept that spans across various domains, each harnessing its unique form of imaginative thought and innovation. In this section, we will review the state of creativity in AI across four major domains where machine creativity is most extensively explored: linguistics, art, science, and problem-solving.
3.1. Linguistic Creativity
@@
The creative aspect of language in linguistics has been discussed since the early days (Chomsky, 1965). Chomsky, in this paper, attributes creativity mainly to the essential property of language to provide means to express many thoughts indefinitely. However, several linguists since Chomsky have argued against using this characterization since it does not align with the everyday definition of creativity (Bergs, 2019; Sampson, 2017; Zawada, 2006). Chomsky’s theory of grammar might generate an infinite number of sentences; it, however, relies on a fixed set of rules, while creativity requires deviation from rules. In this sense, Sampson (2017) suggests distinguishing between F-creativity (fixed) and E-creativity (extending), where F-creativity refers to the Chomskian interpretation of linguistic creativity (a.k.a productivity in morphology) and E-creativity corresponds to the real linguistic innovation such as metaphors, jokes, neologisms, etc. Some recent works have explored the F-creativity of large language models and found that this task is challenging in general and even harder in more morphological complex languages (Ismayilzada et al., 2024a; Anh et al., 2024; Weissweiler et al., 2023). Most past works however have focused on studying the E-creativity of AI systems which we review in the following sections.
3.1.1. Humor
@@
Humor is one of the most common ways in which humans creatively use language to express their ideas and feelings. Early works to model humor focused on hand-crafted linguistic templates and wordplay (Stock and Strapparava, 2005; Taylor and Mazlack, 2004; Raskin and Attardo, 1994). Subsequent works have leveraged language’s lexical and syntactic properties as humor-specific features for humor detection (Liu et al., 2018b; Yang et al., 2015; Zhang and Liu, 2014). The growing interest in computational humor in recent years has resulted in several shared tasks organized by the NLP community (Meaney et al., 2021; Hossain et al., 2020a; Van Hee et al., 2018; Castro et al., 2018; Potash et al., 2017; Miller et al., 2017). Latest works have developed methods based on neural networks and language models to generate and detect humorous content (Ravi et al., 2024; Arora et al., 2022; Peyrard et al., 2021; Annamoradnejad and Zoghi, 2020; Ziser et al., 2020; Amin and Burghardt, 2020; Hossain et al., 2019; Chen and Soo, 2018; Bertero and Fung, 2016), jokes (Horvitz et al., 2024; Tang et al., 2022; Xie et al., 2021; Weller and Seppi, 2019; Ren and Yang, 2017), puns (He et al., 2019; Mittal et al., 2022; Yu et al., 2018), and sarcasm (Chakrabarty et al., 2022c, 2020a). Several datasets have also been proposed to benchmark the humor capacity of LLMs in several languages including English (Horvitz et al., 2024; Jain et al., 2024; Tang et al., 2022; Meaney et al., 2021; Hossain et al., 2020b, 2019; Miller et al., 2017), Chinese (Zhang et al., 2019b), Italian (Buscaldi and Rosso, 2007), Spanish (Castro et al., 2017), Dutch (Winters and Delobelle, 2020) and Russian (Blinov et al., 2019). Computational humor has also been explored in multimodal settings involving images, audio, and video in addition to text (Hessel et al., 2023a; Xie et al., 2023c; Hasan et al., 2019; Bertero and Fung, 2016; Radev et al., 2016; Shahaf et al., 2015). While the latest methods particularly LLMs show an impressive ability to generate and detect humorous content, recent work has also shown that these models still fail to reliably understand humor (Hessel et al., 2023a; Góes et al., 2023a; Kocoń et al., 2023; Borji, 2023a) and generated jokes typically lack diversity (Jentzsch and Kersting, 2023) which has been attributed to training on less diverse humor datasets (Baranov et al., 2023). Creative training frameworks have also been developed to improve the humor generation capabilities of LLMs (Zhong et al., 2023). We refer the reader to Amin and Burghardt (2020) for an in-depth survey on computational humor.
3.1.2. Figurative Language